OpenStudy (anonymous):

The minimum and maximum temperature on a cold day in Lollypop Town can be modeled by the equation below: 2|x − 6| + 14 = 38 What are the minimum and maximum temperatures for this day?

OpenStudy (anonymous):

Answer: Subtract 14 from both sides, 2|x - 6| + 14 - 14 = 38 - 14 2|x - 6| = 24 Divide both sides by 2, [ 2|x - 6| ] / 2 = 24 / 2 |x - 6| = 12 -12 = x - 6 = 12 -12 = x - 6 -12 + 6 = x -6 = x x - 6 = 12 x = 12 + 6 x = 18 x = -6, 18 The minimum T = -6oC The maximum T = 18oC [ x = −6, x = 18 ]